OpenStudy (anonymous):

Please help! how do you write y=x^2-2x-3 in vertex and intercept form using the quadratic formula?

OpenStudy (anonymous):

@ganeshie8 @satellite73 do you guys know how to do this? :o

OpenStudy (anonymous):

I think it is y=-2x-3

OpenStudy (anonymous):

Oops I meant xD y=-2x+1

OpenStudy (anonymous):

is that for vertex form or intercept form? :)

OpenStudy (anonymous):

Oops damn I put that in slope-intercept... Geez I suck at math xD, I don't think I can help you on vertex

OpenStudy (anonymous):

hahah thanks for trying :)

ganeshie8 (ganeshie8):

use the quadratic formula and find the x-intercepts first

OpenStudy (anonymous):

can you help me figure out how to do that? thats where im stuck :(

ganeshie8 (ganeshie8):

sure :) u knw the quadratic formula ?

OpenStudy (anonymous):

sorta :o im just so confused because we've been using so many different forms!

ganeshie8 (ganeshie8):

dont wry, this is going to be very easy :) First, we find x-intercepts using the quadratic formula.

ganeshie8 (ganeshie8):

y=x^2-2x-3 a = ? b = ? c = ?

ganeshie8 (ganeshie8):

\(\large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\)

ganeshie8 (ganeshie8):

you need to figure out a, b, c values. and plug them above to get the x-intercepts

OpenStudy (anonymous):

so a=1 b=2 c=3? what i dont get is that it wants me to put that equation in intercept and vertex form but those have h & k :o

ganeshie8 (ganeshie8):

b = -2 we will get to vertex/intercept forms later. first, get the x-intercepts :)

ganeshie8 (ganeshie8):

put them in the formula above, wat do u get for x = ?

OpenStudy (anonymous):

1 becaue 1^2 is 1? let me do that equation

ganeshie8 (ganeshie8):

a = 1 b = -2 c = -3

ganeshie8 (ganeshie8):

use above, take ur time :)

ganeshie8 (ganeshie8):

\(\large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\) \(\large x = \frac{--2 \pm \sqrt{(-2)^2-4 \times 1\times (-3)}}{2(1)}\)

OpenStudy (anonymous):

i got 1 ± 4 .... is that close? :o

ganeshie8 (ganeshie8):

\(\large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\) \(\large x = \frac{--2 \pm \sqrt{(-2)^2-4 \times 1\times (-3)}}{2(1)}\) \(\large x = \frac{2 \pm \sqrt{4 + 12}}{2}\) \(\large x = \frac{2 \pm \sqrt{16}}{2}\)

ganeshie8 (ganeshie8):

thats close, just a minor mistake u did... check once

ganeshie8 (ganeshie8):

\(\large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\) \(\large x = \frac{--2 \pm \sqrt{(-2)^2-4 \times 1\times (-3)}}{2(1)}\) \(\large x = \frac{2 \pm \sqrt{4 + 12}}{2}\) \(\large x = \frac{2 \pm \sqrt{16}}{2}\) \(\large x = \frac{2 \pm 4}{2}\) \(\large x = 1 \pm 2\)

OpenStudy (anonymous):

ohhh i see what i did wrong! thank you :) so know what do i do?

ganeshie8 (ganeshie8):

that gives, \(x = 1+2 ~~~ or~~~~ x = 1-2\) \(x = 3, -1\)

ganeshie8 (ganeshie8):

we're done wid using quadratic formula. we got the x-intercepts : 3, -1

ganeshie8 (ganeshie8):

so the intercept form would be : y = (x-3)(x+1)

ganeshie8 (ganeshie8):

wat about vertex form ? how to get \(h\), \(k\) ... any ideas ?

OpenStudy (anonymous):

wow that was easier than i thought! thank you :) & no, not too sure :(

ganeshie8 (ganeshie8):

vertec form is also easy, \(h\) lies exactly at the middle of both the intercepts :- \(h \) = \(\large \frac{3 + -1}{2}\)

ganeshie8 (ganeshie8):

\(h \) = \(\large \frac{3-1}{2}\)

ganeshie8 (ganeshie8):

\(h \) = \(\large \frac{2}{2}\)

OpenStudy (anonymous):

so h=1?

ganeshie8 (ganeshie8):

\(h \) = \(\large 1\)

OpenStudy (anonymous):

yay :D

ganeshie8 (ganeshie8):

so vertex form wud be : \(y = (x-1)^2 + k\)

ganeshie8 (ganeshie8):

you still need to find \(k\)

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

okay, so how do we find k now? :)

ganeshie8 (ganeshie8):

put x = h = 1, in given quadratic. it gives u the \(k\) value

ganeshie8 (ganeshie8):

y=x^2-2x-3 put x = 1

ganeshie8 (ganeshie8):

y = 1^2 - 2(1) - 3 = 1 -2 -3 = ?

OpenStudy (anonymous):

-4?

ganeshie8 (ganeshie8):

yup ! thats ur \(k\)

OpenStudy (anonymous):

y=(x−1)^2+4 would be the answer? :)

ganeshie8 (ganeshie8):

who ate the - sign before 4 ?

OpenStudy (anonymous):

i messed up again ._. so it would be y=(x−1)2-4?

ganeshie8 (ganeshie8):

yes :)

OpenStudy (anonymous):

thank you so much!! you're amazing!! :) if it isnt too much too ask could you help me with two more equations similar to that? :)

OpenStudy (anonymous):

i need to figure out how to put y-5=-2(x+1)^2 in standard and intercept form & y=(x+2)(x-3) in standard and vertex form :o

ganeshie8 (ganeshie8):

wats the standard form ?

OpenStudy (anonymous):

y=ax^2+bx+c?

ganeshie8 (ganeshie8):

y = ax^2 + bx + c this is the standard form right ?

ganeshie8 (ganeshie8):

oh yes... thats right !

ganeshie8 (ganeshie8):

y-5=-2(x+1)^2 add 5 both sides y = -2(x+1)^2 + 5

ganeshie8 (ganeshie8):

next, use the formula \((a+b)^2 = a^2 +b^2 + 2ab\)

ganeshie8 (ganeshie8):

expand right side,

OpenStudy (anonymous):

would it be y=-2x+1

ganeshie8 (ganeshie8):

y = -2(x+1)^2 + 5 = -2(x^2 + 1^2 + 2x ) + 5 = -2(x^2 + 1 + 2x) + 5

ganeshie8 (ganeshie8):

simplify :)

OpenStudy (anonymous):

would it be -2x^2-4x+3?

ganeshie8 (ganeshie8):

correct

ganeshie8 (ganeshie8):

standard form : y = -2x^2-4x+3

OpenStudy (anonymous):

yesss :D

ganeshie8 (ganeshie8):

y-5=-2(x+1)^2 to put this guy in intercept form, u need to find its x-intercepts. so put y = 0 0-5=-2(x+1)^2 5/2 = (x+1)^2 \(\pm \sqrt{5/2} = x+1\) \( -1 \pm \sqrt{5/2} = x\)

ganeshie8 (ganeshie8):

those are our x-intercepts

ganeshie8 (ganeshie8):

can u put t he intercept form ?

OpenStudy (anonymous):

the intercept form is y=a(x-p)(x-q)

OpenStudy (anonymous):

so if the equation is −1±5/2, how do i solve the fraction part?